# On the Effectiveness of Large Language Models in Domain-Specific Code Generation

### v2

@@ 4.4. Experimental Setup
Collecting API Knowledge To evaluate code generation with domain knowledge, we extend the domain-specific dataset for RQ1 and RQ2 (Section 3.1) with domain knowledge. We concerning knowledge about certain libraries, including the APIs and their corresponding natural language docstring, organized in a dictionary format. As Table 5 depicts, an API name is a unique identifier of a specific function in a third-party library, a description provides a supplementary and more detailed explanation of its behavior.
@@
Dataset gin prometheus grpc-go Unreal Engine cocos2d-x bgfx
@@
zero-shot
 PolyCoder 8.13 17.05 1.77 12.11 55.36 57.52 0.94 9.87 13.83 23.83 0.76 7.16
 -w/ kg-GPT 9.91 18.16 2.29 12.76 56.87 58.30 1.26 10.20 15.33 24.09 0.78 7.34
 -w/ CoT-PT 9.17 17.41 2.21 12.65 55.97 57.70 1.38 10.15 16.13 24.91 0.91 7.87
@@
 -w/ kg-GPT 13.12 20.44 3.75 14.74 60.41 61.85 1.74 10.98 22.43 27.53 1.20 8.67
 -w/ CoT-PT 12.46 19.77 3.32 13.66 59.91 61.25 1.64 10.96 20.78 25.81 1.17 8.42
finetuning
 PolyCoder 20.62 27.18 5.88 17.11 62.97 63.37 1.57 11.10 27.65 32.67 1.01 11.80
 -w/ CoT-FT 21.12 27.51 7.59 19.13 63.23 63.83 2.26 11.69 29.30 33.99 1.05 11.95
@@
 kg-GPT 45.33 0.36 23.99 0.20 52.64 0.67 15.09 0.09 56.02 0.50 31.02 0.27
Backbone Models We employ PolyCoder (Xu et al., 2022) and StarCoder (Li et al., 2023a) as backbone LLMs, which are well-established open-source models trained on GitHub code. Considering our budgets and available resources, all fine-tuning experiments were only carried out on PolyCoder. PolyCoder is based on GPT-2, which is easy to develop and deploy. The computational cost is affordable by our resources and budgets. More importantly, it demonstrates competitive performance to GPT-3 counterparts (Xu et al., 2022). We employ the released checkpoint of pre-trained PolyCoder-2.7B and StarCoder-15.5B [^7] https://huggingface.co/bigcode/starcoder for all experiments. We implement the API inquirer (kg-GPT) by employing a standard GPT-2 model with $12$ layers and $768$ dimensionality.
@@ 4.5. Results
Domain-finetuning improves the results compared to zero-shot, as it allows the model to learn about domain-specific code most straightforwardly. Still, by injecting the knowledge into training data in a chain-of-thought fashion, the performance is raised to a higher level.
@@ 5.1. Future Directions
First, programming knowledge can extend beyond API documentation. We suggest future research to consider more sources of domain knowledge in the design of prompts and the integration of knowledge. Regarding code syntax knowledge, more comprehensive knowledge of third-party libraries could be helpful, including class descriptions, relationships between classes, and other relevant information. For code semantics knowledge, descriptive and explanatory information from sources like Stack Overflow could be useful.
@@ 6. Conclusion
In this paper, we conduct an empirical study on domain-specific code generation by LLMs. Our study finds that LLMs such as ChatGPT exhibit sub-optimal performance in generating domain-specific code, and adding knowledge prompts about domain libraries improves the performance. To further investigate how to incorporate API knowledge into LLMs, we design three experimental strategies to automatically prompt LLMs, including inquiring an external knowledge GPT, chain-of-thought prompting, and chain-of-though fine-tuning. Our experimental results show the effectiveness of knowledge integration in both zero-shot and fine-tuning settings. We hope our study can inspire future work on improving the code generation capabilities of LLMs for specific domains.

### v3

@@ 4.4. Experimental Setup
Collecting API Knowledge To evaluate code generation with domain knowledge, we extend the domain-specific dataset for RQ1 and RQ2 (Section 3.1) with domain knowledge. We are concerned with knowledge about certain libraries, including the APIs and their corresponding natural language docstring, organized in a dictionary format. As Table 5 depicts, an API name is a unique identifier of a specific function in a third-party library, and an API docstring provides a supplementary and more detailed explanation of its behavior.
@@
Model      gin prometh. grpc-go UnrealEng. cocos2d-x bgfx
@@
PolyCoder (zero-shot) 8.13  17.05 1.77 12.11 55.36 57.52 0.94 9.87 13.83 23.83 0.76 7.16
+In-context learning (Ahmed and Devanbu, 2022) 8.18 16.36 2.01 12.57 55.54 57.11 0.96 0.98 13.27 23.00 0.83 7.24
+DomCoder (kg-GPT) 9.91 18.16 2.29 12.76 56.87 58.30 1.26 10.20 15.33 24.09 0.78 7.34
+DomCoder (CoT-PT) 9.17 17.41 2.21 12.65 55.97 57.70 1.38 10.15 16.13 24.91 0.91 7.87
@@
+DomCoder (CoT-FT) 21.12 27.51 7.59 19.13 63.23 63.83 2.26 11.69 29.30 33.99 1.05 11.95
StarCoder (zero-shot) 8.46 17.55 1.68 12.37 58.69 60.09 0.58 9.00 14.08 23.56 0.80 7.21
+In-context learning (Ahmed and Devanbu, 2022) 7.11 17.32 2.00 12.75 58.12 59.07 1.14 10.76 18.62 26.55 1.12 7.93
+DomCoder (kg-GPT) 13.12 20.44 3.75 14.74 60.41 61.85 1.74 10.98 22.43 27.53 1.20 8.67
+DomCoder (CoT-PT) 12.46 19.77 3.32 13.66 59.91 61.25 1.64 10.96 20.78 25.81 1.17 8.42
CodeLlama (zero-shot) 8.47 18.07 1.79 13.13 57.61 60.39 0.73 10.21 16.76 25.87 0.55 7.72
+In-context learning (Ahmed and Devanbu, 2022) 8.22 18.58 2.69 13.87 59.35 61.49 0.81 10.35 17.90 27.06 0.60 8.47
+DomCoder (kg-GPT) 11.71 19.83 2.62 13.41 55.93 58.24 0.99 10.55 19.58 27.71 0.65 8.66
+DomCoder (CoT-PT) 11.75 19.90 3.07 14.58 56.03 58.69 1.15 11.15 16.39 25.16 0.57 8.21
@@
CodeLlama 21.64 0.06 12.42 0.04 36.92 0.38 9.50 0.07 35.91 0.36 6.73 0.03
kg-GPT 46.48 0.36 27.16 0.20 52.67 0.67 16.48 0.09 57.24 0.50 34.25 0.28
Backbone Models We employ PolyCoder (Xu et al., 2022), StarCoder (Li et al., 2023a), and CodeLlama (Roziere et al., 2023) as backbone LLMs, which are well-established open-source models trained on GitHub code. Considering our computational resources, all fine-tuning experiments were only carried out on PolyCoder. PolyCoder is based on GPT-2, which is easy to develop and deploy. More importantly, it demonstrates competitive performance to GPT-3 counterparts (Xu et al., 2022). We employ the released checkpoint of pre-trained PolyCoder-2.7B, StarCoder-15.5B [^6] https://huggingface.co/bigcode/starcoder, and CodeLlama-7B [^7] https://huggingface.co/codellama/CodeLlama-7b-hf for all experiments. We implement the API inquirer (kg-GPT) by employing a standard GPT-2 model with $12$ layers and $768$ dimensionality.
@@
Comparison Methods We compare our approach with all backbone LLMs in the zero-shot setting, which refers to prompting the pre-trained checkpoint without any further fine-tuning. We also compare DomCoder with a few-shot learning method proposed by Ahmed et al. (Ahmed and Devanbu, 2022), which prepend a few in-context examples in the LLM context. We randomly sample 10 in-context examples that invoke the same library to the target domain. Besides the zero-shot setting, we are also concerned with the fine-tuning strategy of our method. Specifically, we compare DomCoder (CoT-FT) with PolyCoder fine-tuned on the domain-specific dataset.
@@ 4.5. Results
Domain-finetuning improves the results compared to zero-shot, as it allows the model to learn about domain-specific code most straightforwardly. Still, by injecting the knowledge into training data in a chain-of-thought fashion, the performance is raised to a higher level. We note that the improvement that CoT-FT brings to PolyCoder is relatively marginal. Nevertheless, it demonstrates that the proposed chain-of-thought knowledge integration is also useful in the fine-tuning mode despite still having ample room for further improvement.
@@ 4.6. Qualitative Analysis
4.7. Industry Study
To fully assess the effectiveness of DomCoder, we perform an in-house industry study on Tencent Inc, a world-famous IT company. Our study focuses on the domain of messenger Apps (e.g., WeChat, QQ) which are the predominant products of this company. We start by creating a survey about the domain libraries that are mostly adopted by the developers in the group of messenger products. Results show that an internal library called tRPC-go (https://github.com/trpc/trpc) ranked as the most popular one that is utilized by the group. We therefore selected functions that utilize this library and evaluate the performance of DomCoder in completing these functions. To ensure that the selected functions are representative and distinct, we specifically chose functions containing 3 to 7 APIs. Functions with identical API sequences were clustered into groups. These function groups were then sorted based on the number of functions they contained, in descending order. From each class, we randomly selected one function. Additionally, we manually filtered out functions that developers may find challenging to understand. As a result, we obtained a set of 30 test functions for evaluation.
We built DomCoder based on CodeLlama-7B and applied it to complete the selected test functions. We test the kg-GPT strategy which demonstrates the best performance in the quantitative study and is easier to deploy. We measure the performance using two metrics: 1) task-relevance, which means that the generated content satisfies the intention in the function signature, and 2) semantic correctness, which means that the implemented function body is semantically correct. We asked three developers from the relevant team in Tencent to rate the evaluation criteria independently. Conflicts were resolved through discussion until a consensus was reached.
The results are presented in Table 8. We find that DomCoder significantly enhances the quality of generated code. The semantic correctness of the generated code by DomCoder is triple that of the code generated by the original CodeLlama-7B. The results confirm the effectiveness of DomCoder in wider application scenarios and metrics in the industry.
Table 8. Performance of DomCoder in the industry study
Model Task Semantic
Relevance Correctness
CodeLlama-7B (zero-shot) 20/30 6/30
- w/ DomCoder (kg-GPT) 28/30 18/30
@@ 5.1. Future Directions
First, programming knowledge can extend beyond API documentation. We suggest future research to consider more sources of domain knowledge in the design of prompts and the integration of knowledge. Regarding syntax, more comprehensive information about third-party libraries could be helpful, including class descriptions, relationships between classes, and other relevant information. For code semantics, descriptive and explanatory information from sources like Stack Overflow could be useful. Moreover, knowledge about a certain domain such as workflows and code templates could also be useful.
@@ 6. Conclusion
In this paper, we conduct an empirical study on domain-specific code generation by LLMs. Our study finds that LLMs such as ChatGPT exhibit sub-optimal performance in generating domain-specific code, and adding knowledge prompts about domain libraries improves the performance. To further investigate how to incorporate API knowledge into LLMs, we design three experimental strategies to automatically prompt LLMs, including inquiring about an external knowledge GPT, chain-of-thought prompting, and chain-of-though fine-tuning. Our experimental results show the effectiveness of knowledge integration in both zero-shot and fine-tuning settings. We hope our study can inspire future work on improving the code generation capabilities of LLMs for specific domains.